OpenStudy (anonymous):

help!?!?! Explain the difference between using the sine ratio to solve for a missing angle in a right triangle versus using the co-secant ratio. You must use complete sentences and any evidence needed (such as an example) to prove your point of view.

OpenStudy (solomonzelman):

Id get the question but here some thing that might help.  \[\huge\color{red}{Sin~ V= \frac{opp}{hyp} } \] \[\huge\color{red}{Csc~ V= \frac{hyp}{opp} } \] \[\huge\color{green}{Cos~ V= \frac{adj}{hyp} } \] \[\huge\color{green}{Sec~ V= \frac{hyp}{adj} } \] \[\huge\color{blue}{Tan~ V= \frac{opp}{adj} } \] \[\huge\color{blue}{Cot~ V= \frac{adj}{opp} } \]

OpenStudy (anonymous):

@hartnn help plz

OpenStudy (anonymous):

when you know hypotenuse and another side thats opposite to missing angle in a right triangle, you can use 'sin' or 'co-secant' ratios :- sin = opposite/hypotenuse co-secant = 1/sin so, there is no difference. co-secant is just reciprocal of sin. you can use either of them. However you will end up using sine ratio most of the time, as most calculators allow u to get the inverse sin easily without changing settings.

OpenStudy (anonymous):

@beccaboo022a thx

OpenStudy (anonymous):

you wlcme ^_^

OpenStudy (wolf1728):



OpenStudy (anonymous):

@wolf1728 thx

OpenStudy (wolf1728):

u r welcome luka here's a clickable link to the calculator: http://www.1728.org/trigcalc.htm